# thanks: Corresponding author: Sebastian Schulz

### v1

@@ I Introduction
We show that modifications to the problem Hamiltonian itself can greatly improve the performance of iterative QA on noisy hardware. With Learning-Driven Annealing (LDA), we propose a framework to link individual QA runs into a global solution strategy to mitigate hardware constraints such as finite annealing times and integrated control errors [QPU_ICE]. LDA works by analyzing the states sampled from QA to adaptively modify the problem Hamiltonian to the information learned about the energy landscape. This energetically isolates low-energy valleys in the instantaneous energy spectrum, focussing subsequent annealing runs into a promising region of the Hilbert space. As a result, the QPU can reach deeper regions of the energy spectrum that are otherwise inaccessible, enabling NISQ devices to tackle any COP that can be embedded onto their hardware.
@@
We demonstrate the capabilities of LDA by designing a hybrid quantum-classical solver for large-scale spin glasses, that is based on alternating local and global search protocols. Using the D-Wave Advantage 5.4 QPU [QPU_Advantage_5_4] in Jülich, Germany, we benchmark our hybrid solver against leading quantum and classical algorithms, including reverse annealing [IRA_Theory_1], cyclic annealing [Cyclic_annealing_1, Cyclic_annealing_2], simulated annealing using JUPTSA [JUPTSA], Gurobi [Gurobi], Toshiba’s SBM [Toshiba_SBM], VeloxQ [VeloxQ] and D-Wave hybrid [D_Wave_Hybrid], on $5580$-qubit NAT-7 [NAT_7_1, NAT_7_2] spin glasses. The results show that our hybrid solver outperforms all competing strategies in both runtime and lowest energy. LDA is a step towards practical quantum computation that enables today’s quantum annealers to compete with classical solvers.
@@
The remainder of this paper is organized as follows: In Sec. II, we introduce the spin-glass formulation and provide insights into its underlying structures. Section LABEL:sec:qa concerns adiabatic quantum evolutions and discusses its limitations on modern quantum annealers. In Sec. LABEL:sec:lda, we introduce the LDA framework, showcasing how learned information from sampled states can be used to guide the annealing evolution by modifying the problem Hamiltonian. In Sec. LABEL:sec:hybrid_optimization, we demonstrate how LDA can be used in advanced algorithms by designing a hybrid solver for spin glasses based on alternating local and global search protocols. Finally, in Sec. LABEL:sec:benchmarks, we benchmark our hybrid solver against various classical and quantum solvers, followed by a summary of our findings in Sec. LABEL:sec:conclusion.
@@
Figure 1: A $24$-qubit spin-glass with $N_{dom}=3$. (a) Spin-glass graph with the biases $h_{i}$ as nodes and couplers $J_{ij}$ as edges. Dashed edges indicate periodic boundary conditions. The colored rectangles represent the spin domains: $\mathcal{D}_{1}$ (green), $\mathcal{D}_{2}$ (purple), and $\mathcal{D}_{3}$ (orange). Each domain has two distinct minima: $\alpha$ and $\overline{\alpha}$ ($\mathcal{D}_{1}$), $\beta$ and $\overline{\beta}$ ($\mathcal{D}_{2}$), $\gamma$ and $\overline{\gamma}$ ($\mathcal{D}_{3}$). Panels (a.1-3) show the eigenenergies $E$ of the corresponding domains as a function of the Hamming distance $d_{ham}$ to the respective ground state $\alpha,\beta,\gamma$. (b) Hypercube representation of the Hilbert space, with each axis representing the Hamming distance $d_{ham}$ from the ground state ($\alpha,\beta,\gamma$) of the respective spin domains $\mathcal{D}_{k}$. Color indicates the lowest eigenenergy at each point. The corners of the hypercube correspond to the $82^{N_{dom}}=$ 8 local minima of the spin glass. (c) Energy landscape of the spin glass obtained by unfolding the hypercube facets. Each point is projected onto the closest facet (smallest Hamming distance), with the blue surface representing the lowest eigenenergies on these facets. The eight local minima are marked by colored diamonds at the facets’ corners, with the states surrounding it being the energy valleys. Additionally, black dots indicate $2000$ samples from a quantum annealing simulation with annealing time $T=10$, with the probability density shown on the lower orange plane. Panels (c.1-3) highlight the specific biases $h_{i}$ and couplers $J_{ij}$ satisfied by three of the eight local minima, with the unsatisfied terms set to 00 (white). The color scheme matches that of panel (a).
II Spin-glass problems
@@
Spin glasses (SG) are magnetic systems of quenched disorder, imposing conflicts between interacting magnetic moments [SG_Theory]. In the Edwards-Anderson model of SG [SG_Theory_EA], $N$ Ising spins $\sigma^{z}_{i}$, $i\in\{0,\,\dots\,,\,N-1\}$, are placed on the sites of a regular graph, with the Hamiltonian defined as
@@
Here $J_{ij}$ with $i<j$ denotes the coupling strength between spins $i$ and $j$, and $h_{i}$ represents a locally applied external magnetic field bias on spin $i$.
We consider both triangular lattices with periodic boundary conditions (see Fig. 1a) and the Pegasus graph used by the D-Wave Advantage 5.4 quantum annealer [QPU_Advantage_5_4]. In the following, we choose |$\lvert$|

### v2

@@ 1 Introduction
We show that systematically learned modifications to the problem Hamiltonian itself can greatly improve the performance of iterative QA on noisy hardware. With Learning-Driven Annealing (LDA), we propose a framework to link individual QA runs into a global solution strategy to mitigate hardware constraints such as finite annealing times and integrated control errors [QPU_ICE]. LDA works by analyzing the states sampled from QA to adaptively modify the problem Hamiltonian to the information learned about the energy landscape. This energetically isolates low-energy valleys in the instantaneous energy spectrum, focussing subsequent annealing runs into a promising region of the Hilbert space. As a result, the QPU can reach deeper regions of the energy spectrum that are otherwise inaccessible. While we do not claim a quantum advantage per se, LDA enables current NISQ devices to tackle COPs more efficiently as long as they can be embedded onto their hardware.
@@
We demonstrate the capabilities of LDA by designing a hybrid quantum-classical solver for large-scale spin glasses, that is based on alternating local and global search protocols. Using the D-Wave Advantage1 5.4 QPU [QPU_Advantage_5_4] in Jülich, Germany, we benchmark our hybrid solver against leading quantum and classical algorithms, including reverse annealing [IRA_Theory_1], cyclic annealing [Cyclic_annealing_1, Cyclic_annealing_2], simulated annealing using JUPTSA [JUPTSA], Gurobi [Gurobi], Toshiba’s SBM [Toshiba_SBM], VeloxQ [VeloxQ, VeloxQ2] and D-Wave hybrid [D_Wave_Hybrid], on $5580$-qubit NAT-7 [NAT_7_1, NAT_7_2] spin glasses. The results show that our hybrid solver outperforms all competing strategies in both runtime and lowest energy. LDA is a step towards practical quantum computation that enables today’s quantum annealers to compete with classical solvers.
@@
The remainder of this paper is organized as follows: In Sec. 2, we introduce the spin-glass formulation and provide insights into its underlying structures. Section LABEL:sec:qa concerns adiabatic quantum evolutions and discusses its limitations on modern quantum annealers. In Sec. LABEL:sec:lda, we introduce the LDA framework, showcasing how learned information from sampled states can be used to guide the annealing evolution by modifying the problem Hamiltonian. In Sec. LABEL:sec:hybrid_optimization, we demonstrate how LDA can be used in advanced algorithms by designing a hybrid solver for spin glasses based on alternating local and global search protocols. Finally, in Sec. LABEL:sec:benchmarks, we benchmark our hybrid solver against various classical and quantum solvers, followed by a summary of our findings in Sec. LABEL:sec:conclusion.
@@
Figure 1: A $24$-qubit spin-glass with $N_{dom}=3$. (a) Spin-glass graph with the biases $h_{i}$ as nodes and couplers $J_{ij}$ as edges. Dashed edges indicate periodic boundary conditions. The colored rectangles represent the spin domains: $\mathcal{D}_{1}$ (green), $\mathcal{D}_{2}$ (purple), and $\mathcal{D}_{3}$ (orange). Each domain has two distinct minima: $\alpha$ and $\overline{\alpha}$ ($\mathcal{D}_{1}$), $\beta$ and $\overline{\beta}$ ($\mathcal{D}_{2}$), $\gamma$ and $\overline{\gamma}$ ($\mathcal{D}_{3}$). Panels (a.1-3) show the eigenenergies $E$ of the corresponding domains as a function of the Hamming distance $d_{ham}$ to the respective ground state $\alpha,\beta,\gamma$. (b) Hypercube representation of the Hilbert space, with each axis representing the Hamming distance $d_{ham}$ from the ground state ($\alpha,\beta,\gamma$) of the respective spin domains $\mathcal{D}_{k}$. Color indicates the lowest eigenenergy at each point. The corners of the hypercube correspond to the $2^{N_{dom}}=8$ local minima of the spin glass. (c) Energy landscape of the spin glass obtained by unfolding the hypercube facets. Each point is projected onto the closest facet (smallest Hamming distance), with the blue surface representing the lowest eigenenergies on these facets. The eight local minima are marked by colored diamonds at the facets’ corners, with the states surrounding it being the energy valleys. Additionally, black dots indicate $2000$ samples from a quantum annealing simulation with annealing time $T=10$, with the probability density shown on the lower orange plane. Panels (c.1-3) highlight the specific biases $h_{i}$ and couplers $J_{ij}$ satisfied by three of the eight local minima, with the unsatisfied terms set to 0 (white). The color scheme matches that of panel (a).
2 Spin-glass problems
@@
Spin glasses (SG) are magnetic systems of quenched disorder, imposing conflicts between interacting magnetic moments [SG_Theory]. In the Edwards-Anderson model of SG [SG_Theory_EA], NN Ising spins $\sigma^{z}_{i}$, $i\in\{0,\,\dots\,,\,N-1\}$, are placed on the sites of a regular graph, with the Hamiltonian defined as
@@
Here $J_{ij}$ with $i<j$ denotes the coupling strength between spins $i$ and $j$, and $h_{i}$ represents a locally applied external magnetic field bias on spin ii.
We consider both triangular lattices with periodic boundary conditions (see Fig. 1a) and the Pegasus graph used by the D-Wave Advantage1 5.4 quantum annealer [QPU_Advantage_5_4]. In the following, we choose $\lvert$